# An Online Learning Approach to Prompt-based Selection of Generative Models

### v2

@@ 3.1 CLIPScore
and note that $\cos(\bm{v}_{x},\bm{c}_{y})=\langle\bm{v}_{x},\bm{c}_{y}\rangle$ as we operate under the normalized embeddings. Further, for a video $X:=\{x^{(l)}\}_{l=1}^{L}$ consisting of $L$ frames, where $x^{(l)}$ is the $l$-th frame, the score is the averaged frame CLIPScore, that is,
$\textup{CLIPScore}^{\textup{T2V}}(y,X):=\frac{1}{L}\sum_{l=1}^{L}\textup{CLIPScore}^{\textup{T2I}}(y,x^{(l)}).$ (2)
3.2 Kernel Methods and Random Fourier Features
Let $\phi:{\mathbb{R}}^{d}\to{\mathcal{H}}$ denote a mapping from the primal space ${\mathbb{R}}^{d}$ to the (possibly infinite-dimensional) associated reproducing kernel Hilbert space (RKHS) ${\mathcal{H}}$. The corresponding kernel function is defined by $k(y,y^{\prime}):=(\phi(y))^{\top}\phi(y^{\prime})$ for any $y,y^{\prime}\in{\mathbb{R}}^{d}$. The kernel function $k$ is positive definite if $\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}k(y_{i},y_{j})\geq 0$ for any $n\in{\mathbb{N}}_{+}$, $y_{1},\cdots,y_{n}\in{\mathbb{R}}^{d}$, and $c_{1},\cdots,c_{n}\in{\mathbb{R}}$. Further, a positive definite kernel function is shift invariant if $k(y,y^{\prime}):=k(y-y^{\prime})$ for any $y,y^{\prime}\in{\mathbb{R}}^{d}$. An example is the radial basis function (RBF) kernel:
$k_{\textup{RBF}}(y,y^{\prime})=\exp(-\|y-y^{\prime}\|_{2}^{2}/(2\sigma^{2})),$
where $\sigma>0$.
Kernel ridge regression (KRR). Given empirical data $(y_{1},s_{1}),\cdots,(y_{n},s_{n})$, where $\{y_{i}\in{\mathbb{R}}^{d}\}_{i=1}^{n}$ are dependent variables and $\{s_{i}\in{\mathbb{R}}\}_{i=1}^{n}$ are target variables, respectively, the kernel method assumes the existence of $w^{\star}\in{\mathcal{H}}$ such that $\mathbb{E}[s_{i}|y_{i}]=(\phi(y))^{\top}w^{\star}$ for any $i=1,\cdots,n$. Let $\alpha\geq 0$ denote a regularization parameter. KRR constructs the estimator
$\widehat{s}_{\textup{KRR}}(y):=k_{y}^{\top}(K+\alpha I_{n})^{-1}v$
for any $y\in{\mathbb{R}}^{d}$, where $K=[k(y_{i},y_{j})]_{i,j=1}^{n}\in{\mathbb{R}}^{n\times n}$ is the kernel matrix, $v:=[s_{1},\cdots,s_{n}]^{\top}\in{\mathbb{R}}^{n}$, and $k_{y}=[k(y_{1},y),\cdots,k(y_{n},y)]^{\top}\in{\mathbb{R}}^{n}$. The estimator can be interpreted as first estimating $w^{\star}$ by ridge regression
$\widehat{w}:=\operatorname*{arg\,min}_{w\in{\mathcal{H}}}\sum_{i=1}^{n}((\phi(y_{i}))^{\top}w-s_{i})^{2}+\alpha\|w\|,$
Random Fourier features (RFF). One problem of KRR is that it scales poorly with the size $n$ of the empirical data, i.e., computing the KRR estimator generally requires $\Omega(n^{3})$ time and $\Omega(n^{2})$ memory. To address this problem, Rahimi and Recht, 2007b propose to scale up kernel methods by RFF sampling. Specifically, the Bochner’s Theorem (Rudin,, 2017) implies that for any (properly scaled) shift-invariant kernel $k(y,y^{\prime})=k(y-y^{\prime})$, there exists a distribution $p\in\Delta({\mathbb{R}}^{d})$ such that $k(y,y^{\prime})=\mathbb{E}_{w\sim p}[e^{iw^{\top}(y-y^{\prime})}],$ where $e^{i\theta}:=\cos\theta+i\cdot\sin\theta$ for any $\theta\in{\mathbb{R}}$ and $i$ is the imaginary unit. Therefore, the idea of RFF is to sample $w_{1},\cdots,w_{s}\sim p$ and approximate $k(y,y^{\prime})$ by the empirical mean $s^{-1}\sum_{j=1}^{s}e^{iw_{j}^{\top}(y-y^{\prime})}$ to within $\epsilon$ with only $s=O(d\epsilon^{-2}\log(1/\epsilon^{2}))$. Since the kernel $k(\cdot)$ is real, we can replace the complex exponentials with cosines and define
$\varphi(y):=\sqrt{\frac{2}{s}}\cdot[\cos(w_{1}^{\top}y+b_{1}),\cdots,\cos(w_{s}^{\top}y+b_{s})]^{\top},$ (3)
[…]
$\widetilde{s}_{\textup{KRR}}(y):=(\widetilde{\Phi}^{*}\widetilde{\Phi}+\alpha I_{s})^{-1}\widetilde{\Phi}^{*}v,$
where $\Phi:=[\varphi(y_{i})^{\top}]_{i=1}^{n}\in{\mathbb{C}}^{n\times s}$, can be computed in $O(ns^{2})$ time and $O(ns)$ memory, giving substantial computational savings if $s\ll n$ (Avron et al.,, 2017). For the RBF kernel, the distribution $p_{\textup{RBF}}$ is the multivariate Gaussian ${\mathcal{N}}(0,\sigma^{-2}\cdot I_{d})$.
@@ 4 Prompt-Based Selection as Contextual Bandits
Protocol 1 Online Prompt-based Selection of Generative Models
@@
5:     Algorithm ${\mathcal{A}}$ picks model $g_{t}\sim{\mathcal{A}}(\cdot|{\mathcal{D}},y_{t})$ and samples an answer $x_{t}\sim P_{g_{t}}(\cdot|y_{t})$.
@@ 5 An Optimism-based Approach for Prompt-based Selection
We note that Assumption 1 differs from the standard settings of linear and kernelized contextual bandit (CB) (Chu et al.,, 2011; Valko et al.,, 2013; Zenati et al.,, 2022). Specifically, in these settings, a different context variable is observed at each arm, and a single weight vector is applied uniformly across all arms. In the prompt-based model selection setting, the text prompt (i.e., the context variable) is shared among arms (i.e., the generative models) and can result in different performance scores across the models, which cannot be captured by the standard kernelized CB approach.
Remark 2 (Kenelized scores).
Note that the relationship between the prompt vector and score can be non-linear and dependent on the choice of the generative model. For instance, the generated images often vary across the prompts and different T2I models, which will affect the resulting CLIPScore (1).
5.1 The PAK-UCB Algorithm
In this section, we present PAK-UCB in Algorithm 2, an online learning approach to prompt-based model selection. For each incoming prompt, PAK-UCB first estimates the performance scores via kernel ridge regression (KRR) and then picks the model with the highest estimated score. Unlike standard LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) algorithms, which train a single reward function shared across all arms, PAK-UCB learns arm-specific functions to predict the score of each model.
@@ 5.1 The PAK-UCB Algorithm
5:     Compute $(\widehat{\mu}_{g},\widehat{\sigma}_{g})\leftarrow\textsc{Compute\_UCB}({\mathcal{D}},y_{t},\Psi_{g})$ for each $g\in{\mathcal{G}}$.
6:     Pick model $g_{t}\leftarrow\operatorname*{arg\,max}_{g\in{\mathcal{G}}}\{\widehat{s}_{g}\}$, where $\widehat{s}_{g}\leftarrow\widehat{\mu}_{g}+(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}$.
7:     Sample an answer $x_{t}\sim P_{g_{t}}(\cdot|y_{t})$ and compute the score $s_{t}\leftarrow s(y_{t},x_{t})$.
@@
The key component in PAK-UCB is the function Compute_UCB (lines 8-17), which outputs both the KRR estimator $\widehat{\mu}_{g}$ and an uncertainty quantifier $\widehat{\sigma}_{g}$. As the weight vector $w_{g}^{\star}$ can vary across the arms, PAK-UCB constructs the KRR dataset using prompt-score pairs from iterations where model $g$ is chosen, with the corresponding indices stored in the set $\Psi_{g}$ (line 6). The estimated score is then computed by $\widehat{s}_{g}=\widehat{\mu}_{g}+(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g}$ (line 4), which is initially set to +$\infty$+ ∞ to ensure each model is picked at least once (lines 9-10). Particularly, the following lemma shows that under some conditions, $\widehat{s}_{g}$ is an optimistic estimation of $s_{g}(y_{t})$ with high probability. The detailed proof can be found in Appendix C.1.
Lemma 1 (Optimism).
We show that a variant of PAK-UCB attains a regret of $\widetilde{O}(\sqrt{GT})$. The formal statement and the proof can be found in Appendix A.
@@
Under the same conditions in Lemma 1, with probability of at least $1-\delta$, a variant of Algorithm 2 attains a regret of $\widetilde{O}(\sqrt{GT})$.
@@ 5.2 PAK-UCB with Random Fourier Features
The PAK-UCB solves a KRR for each model at an iteration to estimate the scores, which can be expensive in both computation and memory for a large number of iterations. To address this problem, we leverage the random Fourier features (RFF) sampling (Rahimi and Recht, 2007b, ) for positive definite shift-invariant kernels. At a high level, RFF maps the input data, e.g., the prompt (vector) in our setting, to a randomized low-dimensional feature space and then applies fast linear methods to solve the regression problem. Particularly, the inner product between these projected randomized features is an unbiased estimation of the kernel value.
We present the RFF-UCB algorithm, which is a variant of PAK-UCB with random features. RFF-UCB leverages an RFF-based approach to compute the mean and uncertainty quantifier in line 3 of Algorithm 2, which we present in Algorithm 3. Upon receiving the regression dataset consisting of prompt-score pairs, Compute_UCB_RFF first projects each $d$-dimensional prompt vector to a randomized $s$-dimensional feature space according to Equation (3) (lines 5-6) and then solves a linear ridge regression to estimate the mean and uncertainty (lines 8-9). To derive statistical guarantees, the number of features varies according to both the input data and error thresholds, which we will specify in Appendix B.1. In practice, we find that a size around 50 can attain satisfactory empirical performance. To see why RFF can reduce the computation, note that the size of the (regularized) Gram matrix $(\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}+\alpha I)$ in line 8 is fixed to be $s$ in the whole process, while the size of $(K+\alpha I)$ in line 13 of Algorithm 2 scales with $|\Psi_{g}|$ and can grow linearly over iterations. Particularly, the following lemma shows that Compute_UCB_RFF can reduce the time and space by an order of $O(t^{2})$ and $O(t)$, respectively.
1:the Fourier transform $p$ of a positive definite shift-invariant kernel $k(y,y^{\prime})=k(y-y^{\prime})$, error thresholds $\operatorname{\epsilon_{\textup{RFF}}},\operatorname{\Delta_{\textup{RFF}}}>0$, regularization and exploration parameters $\alpha,\eta\geq 0$
2:number of features $s$, bonus terms ${\mathcal{B}}_{g,1}$ and ${\mathcal{B}}_{g,2}$
3:function Compute_UCB_RFF(${\mathcal{D}},y,\Psi_{g}$)
4:     if $\Psi_{g}$ is empty then
6:     else
[…]
8:         Define mapping $\varphi(y^{\prime})\leftarrow\sqrt{\frac{2}{s}}\cdot[\cos(w_{1}^{\top}y+b_{1}),\cdots,\cos(w_{s}^{\top}y+b_{s})]^{\top}$ for any $y^{\prime}\in{\mathbb{R}}^{d}$.
10:         $\widetilde{\mu}_{g}\leftarrow(\varphi(y))^{\top}(\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}+\alpha I)^{-1}\widetilde{\Phi}_{g}^{\top}v+{\mathcal{B}}_{g,1}$.
11:         $\widetilde{\sigma}_{g}\leftarrow\alpha^{-\frac{1}{2}}(1-(\varphi(y))^{\top}(\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}+\alpha I)^{-1}\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}(\varphi(y)))^{\frac{1}{2}}+{\mathcal{B}}_{g,2}$.
12:     end if
13:     return $(\widetilde{\mu}_{g},\widetilde{\sigma}_{g})$.
14:end function
Algorithm 3 Compute UCB with Random Fourier Features
@@
At any iteration $t\in[T]$, Compute_UCB (lines 8-17 of Algorithm 2) requires $O(t^{3}/G^{2})$ time and $O(t^{2}/G)$ space, while Compute_UCB_RFF with random features of size $s\in{\mathbb{N}}_{+}$ (Algorithm 3) requires $O(ts^{2})$ time and $O(ts)$ space, where $G$ is the number of generators. See Appendix B.5 for details.
It can be shown that the implementation of PAK-UCB with RFF attains the exact same regret guarantees for adaptively selected feature sizes. The formal statement and the proof can be found in Appendix B.2.
Theorem 2 (Regret when using RFF, informal).
Under the same conditions in Theorem 4, a variant of RFF-UCB attains a regret of $\widetilde{O}(\sqrt{GT})$.
@@ 6 Numerical Results
In this section, we present numerical results for the proposed 1) PAK-UCB-poly3 algorithm (PAK-UCB using a polynomial kernel with degree 3, i.e., $k_{\text{poly3}}(x_{1},x_{2})=(1+x_{1}^{\top}x_{2})^{3}$), and 2) the RFF-UCB algorithm with the RBF kernel. Our primary focus is on prompt-based selection among standard text-to-image (T2I) models: Stable Diffusion v1.5111https://huggingface.co/docs/diffusers/en/api/pipelines/stable_diffusion/text2img, PixArt-$\alpha$-XL-2-512x512222https://huggingface.co/PixArt-alpha/PixArt-XL-2-512x512, UniDiffuser [^3] https://github.com/thu-ml/unidiffuser, and DeepFloyd IF-I-M-v1.0.444https://github.com/deep-floyd/IF. Additionally, we conduct experiments on image captioning (image-to-text) task and video data under synthetic setups. Detailed implementation of the algorithms and additional results can be found in Appendix D.
6.1 Experimental Details
@@
Baselines. We compare PAK-UCB-poly3 and RFF-UCB with five baselines, including 1) LinUCB and KernelUCB-poly3: standard LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) using a polynomial kernel with degree 3, where the context variable of each arm is given by the concatenation of the CLIP embedding of the text prompt and its one-hot indicator vector, 2) One-arm Oracle: always picking the model with the maximum averaged CLIPScore, 3) PAK-UCB-lin: PAK-UCB with linear kernel, i.e., $k_{\text{lin}}(x_{1},x_{2})=x_{1}^{\top}x_{2}$, which does not incorporate non-linearity in score estimation, and 4) Naive-KRR: PAK-UCB-poly3 without exploration, which selects the model with the highest estimated mean conditioned to the prompt.
@@
Performance metrics. For each experiment, we report two performance metrics: (i) outscore-the-best (O2B): the difference between the CLIPScore attained by the algorithm and the highest average CLIPScore attained by any single model, and (ii) optimal-pick-ratio (OPR): the overall ratio that the algorithm picks the best generator conditioned to the prompt type.
6.2 Summary

### v3

@@ 3.2 CLIPScore for Evaluating Text-to-Image Models
where $\cos(\bm{v}_{x},\bm{c}_{y})=\langle\bm{v}_{x},\bm{c}_{y}\rangle$ for the $\ell_{2}$-normalized CLIP-embedded $\bm{v}_{x},\bm{c}_{y}$. Extending the definition to (text,video) pairs, for a video $X:=\{x^{(l)}\}_{l=1}^{L}$ consisting of $L$ frames, where $x^{(l)}$ is the $l$-th frame, the score is the averaged frame-level CLIPScore:
$\textup{CLIPScore}^{\textup{T2V}}(y,X):=\frac{1}{L}\sum_{l=1}^{L}\textup{CLIPScore}^{\textup{T2I}}(y,x^{(l)}).$
@@ 4 Prompt-Based Selection as Contextual Bandits
Protocol 1 Online Prompt-based Selection of LLMs and Prompt-guided Generative Models
@@
5:     Algorithm ${\mathcal{A}}$ picks model $g_{t}\sim{\mathcal{A}}(\cdot|{\mathcal{D}},y_{t})$ and samples response $x_{t}\sim P_{g_{t}}(\cdot|y_{t})$.
@@ 5 An Optimism-based Approach for Prompt-based Selection
We note that Assumption 1 differs from the standard settings of linear and kernelized contextual bandit (CB) (Chu et al.,, 2011; Valko et al.,, 2013; Zenati et al.,, 2022). Specifically, in the settings of these references, a potentially different context variable is considered for each arm, and the same weight vector is applied across all arms. In the prompt-based model selection setting, the text prompt (i.e., the context variable) is shared among arms (i.e., the generative models) and can result in different performance scores across the models, which cannot be captured by the standard kernelized CB approach. On the other hand, we note that the existing arm-specific bandit algorithms in (Li et al.,, 2010; Xu et al.,, 2020) are designed considering a linear pay-off structure, which is generally different from the kernel-based formulation in our work.
@@
5:     Compute $(\widehat{\mu}_{g},\widehat{\sigma}_{g})\leftarrow\textsc{Compute\_UCB}({\mathcal{D}},y_{t},\Psi_{g})$ and set $\widehat{s}_{g}\leftarrow\widehat{\mu}_{g}+\eta\widehat{\sigma}_{g}$ for each $g\in{\mathcal{G}}$.
6:     Pick model $g_{t}\leftarrow\operatorname*{arg\,max}_{g\in{\mathcal{G}}}\{\widehat{s}_{g}\}$.
7:     Sample $x_{t}\sim P_{g_{t}}(\cdot|y_{t})$ and receive score $s_{t}$.
@@
5.1 The PAK-UCB Algorithm
In this section, we present PAK-UCB in Algorithm 2, an online learning approach to prompt-based model selection. For each incoming prompt, PAK-UCB first estimates the performance scores via kernel ridge regression (KRR) and then picks the model with the highest estimated score. Unlike conventional formulations of LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) algorithms, which learn a single reward function shared across all arms, PAK-UCB learns arm-specific functions to predict the score of each model.
The key component in PAK-UCB is the function Compute_UCB (lines 8-17), which outputs both the KRR estimator $\widehat{\mu}_{g}$ and an uncertainty quantifier $\widehat{\sigma}_{g}$. As the weight vector $w_{g}^{\star}$ can vary across the arms, PAK-UCB constructs the KRR dataset using prompt-score pairs from iterations where model $g$ is chosen, with the corresponding indices stored in the set $\Psi_{g}$ (line 6). The estimated score is then computed by $\widehat{s}_{g}=\widehat{\mu}_{g}+\eta\widehat{\sigma}_{g}$ (line 4), which is initially set to +$\infty$+ ∞ to ensure each model is picked at least once (lines 9-10). To provide theoretical justification for our method, we show that a variant of PAK-UCB, which is illustrated in Algorithm 3 in the Appendix, can attain a squared-root regret bound. The formal statement and the proof can be found in Appendix A.
@@ 5.1 The PAK-UCB Algorithm
With probability of at least $1-\delta$, the regret of running a variant of PAK-UCB (stated in Algorithm 3 in the Appendix) for $T$ iterations is bounded by $\widetilde{O}(\sqrt{GT})$.
@@ 5.2 PAK-UCB with Random Fourier Features
The PAK-UCB solves a KRR for each model at an iteration to estimate the scores, which can be expensive in both computation and memory for a large number of iterations. To address this problem, we leverage the random Fourier features (RFF) sampling (Rahimi and Recht, 2007b, ) for shift-invariant kernel functions, e.g., the RBF (Gaussian) kernel. At a high level, RFF maps the input data, e.g., the prompt (vector) in our setting, to a randomized low-dimensional feature space and then learns a linear model in the resulting random space by solving a standard linear regression problem. Note that following the design of RFF, the inner product of the projected randomized features will be an unbiased estimation of the original kernel inner product.
We present RFF-UCB as an RFF implementation of PAK-UCB (Algorithm 4 in the Appendix). Particularly, RFF-UCB leverages an RFF-based approach to compute the mean and uncertainty quantifier, which is stated in Algorithm 5 in the Appendix. Upon receiving the regression dataset consisting of prompt-score pairs, Compute_UCB_RFF first projects each $d$-dimensional prompt vector to a randomized $2D$-dimensional feature space according to Equation (1) and then solves a linear ridge regression to estimate the mean and uncertainty. To see why RFF can reduce the computation, note that the size of the (regularized) Gram matrix $(\widetilde{\Phi}_{g}^{\top}\widetilde{\Phi}_{g}+\alpha I_{2D})$ in Line 8 is fixed to be $2D$ in the whole process, while the size of $(K+\alpha I)$ in line 13 of Algorithm 2 scales with $|\Psi_{g}|$ and can grow linearly over iterations. Particularly, the following lemma shows that Compute_UCB_RFF can reduce the time and space by an order of $O(t^{2})$ and $O(t)$, respectively.
@@
At any iteration $t\in[T]$, Compute_UCB (Lines 8-17 of Algorithm 2) requires $O(t^{3}/G^{2})$ time and $O(t^{2}/G)$ space, while Compute_UCB_RFF with random features of size $s\in{\mathbb{N}}_{+}$ (Algorithm 5) requires $O(tD^{2})$ time and $O(tD)$ space, where $G$ is the number of generators. See Appendix B.1 for details.
To provide theoretical justification to the RFF approach, we show that the implementation of PAK-UCB with RFF can attain the exact same regret bound with a carefully chosen feature sizes. The formal statement and the proof can be found in Appendix B.1.
Theorem 2 (Regret of RFF implementation, informal).
With probability at least $1-\delta$, the regret of running a variant of RFF-UCB (stated in Algorithm 6) for $T$ iterations is bounded by $\widetilde{O}(\sqrt{GT})$.
@@ 6 Numerical Results
In this section, we present numerical results for the proposed 1) PAK-UCB-poly3: PAK-UCB using a polynomial kernel with degree 3, i.e., $k_{\text{poly3}}^{\gamma}(x_{1},x_{2})=(1+\gamma\cdot x_{1}^{\top}x_{2})^{3}$ and 2) RFF-UCB-RBF: PAK-UCB with RFF implementation of the RBF kernel, i.e., $k_{\textup{RBF}}^{\sigma}(x_{1},x_{2})=\exp(-\frac{1}{2\sigma^{2}}\|x_{1}-x_{2}\|^{2})$. Implementation details and the choice for hyperparameters can be found in Appendix C.4. Our primary focus is on prompt-based selection among standard text-to-image (T2I) models: Stable Diffusion v1.5333https://huggingface.co/docs/diffusers/en/api/pipelines/stable_diffusion/text2img, PixArt-$\alpha$-XL-2-512x512444https://huggingface.co/PixArt-alpha/PixArt-XL-2-512x512, UniDiffuser [^5] https://github.com/thu-ml/unidiffuser, and DeepFloyd IF-I-M-v1.0.666https://github.com/deep-floyd/IF. For the LLM experiments, we provide numerical results for prompt-based selection of the following large language models (LLMs): Gemini-2.5-Flash-preview, o3-mini, Deepseek-Chat, and Qwen-Plus. In the Appendix, we report additional results on image captioning (image-to-text) task and video data under synthetic setups. Additional results can be found in Appendix C.
@@
Baselines. We compare PAK-UCB-poly3 and RFF-UCB-RBF with three baselines, including 1) KernelUCB-poly3: standard KernelUCB (Valko et al.,, 2013) using the $k_{\textup{poly}}^{1.0}$ kernel, 2) One-arm Oracle: always picking the model with the maximum averaged score, and 3) Random: selecting an model uniformly randomly. In the Appendix, we report results for three additional baselines: 4) PAK-UCB-lin: PAK-UCB with linear kernel, i.e., $k_{\text{lin}}(x_{1},x_{2})=x_{1}^{\top}x_{2}$, which does not incorporate non-linearity in score estimation, 5) LinUCB: standard LinUCB (Chu et al.,, 2011), and 6) Naive-KRR: PAK-UCB-poly3 without exploration, i.e., parameter $\eta=0$, which selects the model with the highest estimated mean conditioned to the prompt.
@@
Performance metrics. For each experiment, we report two performance metrics: (i) outscore-the-best (OtB): the difference between the CLIPScore attained by the algorithm and the highest average CLIPScore attained by any single model, and (ii) optimal-pick-ratio (OPR): the overall ratio that the algorithm picks the best generator conditioned to the prompt. Specifically, to determine the best model for each specific prompt, we generate five responses from every model (to that prompt), and the best model in the OPR calculation is the model with the highest averaged score over the 5 runs. Then, OPR is defined as the fraction that the algorithm’s selected model matches the best model.
6.1 Summary of the Numerical Results